OpenStudy (anonymous):

You have a mortgage of $110,000 with a term of 30 years at an APR of 9%. Compute your monthly payment.

OpenStudy (ac3):

you have the formula for this?

OpenStudy (anonymous):

Yeah, I have all the formulas I just don't know which one to use

OpenStudy (ac3):

oh, i can help you understand which to use. Post them all up for me please.

OpenStudy (ac3):

I don't remember them it's been a while.

OpenStudy (anonymous):

\[A= P( 1+ \frac{ R }{ M } ) ^{Mt}\]

OpenStudy (ac3):

I think that's the one we want.

OpenStudy (ac3):

What' A stand for? i know P is principal, R is rate, m is months and t is time in years

OpenStudy (ac3):

m depends if it's quarterly monthly continously etc.

OpenStudy (anonymous):

I think so too. So far I have \[A=110,000 \left( 1+\frac{ .09 }{ ? } \right) \]

OpenStudy (anonymous):

I don't know what to insert under .09

OpenStudy (ac3):

so since your doing in monthly M=12. because it's 12 months in a year

OpenStudy (ac3):

and that's the amount of times that the rate will be applied

OpenStudy (anonymous):

Ahhhhh okay

OpenStudy (ac3):

m isn't always months. if it said what's your quarterly payment then you would use 4

OpenStudy (ac3):

because you would apply the rate 4 times in a year.

OpenStudy (ac3):

make sense?

OpenStudy (anonymous):

So let me type out the equation fully

OpenStudy (anonymous):

\[A=110,000 \left( 1+\frac{ 0.9 }{ 12 } \right)^{12(30)}\]

OpenStudy (anonymous):

is that correct?

OpenStudy (ac3):

well done. alright so just to make sure you understand it

OpenStudy (ac3):

what if it said what is your semi-annual payment?

OpenStudy (ac3):

what would M be there?

OpenStudy (anonymous):

2

OpenStudy (ac3):

that's right you got it.

OpenStudy (ac3):

if it says continously you have to use the other equation

OpenStudy (anonymous):

I have an example like that in my notes, haha. Thank you very much!!

OpenStudy (ac3):

\[P=e ^{rt}\]

OpenStudy (ac3):

that's the one if it says "compounded continuously" or something like that

OpenStudy (ac3):

Yw

OpenStudy (anonymous):

Got it!

OpenStudy (ac3):

@ikileyxx wait what did you get for your final answer?

OpenStudy (ac3):

@ikileyxx i think we used the wrong formula

OpenStudy (anonymous):

oh gosh haha uh, $125,558 and monthly payment of $349

OpenStudy (ac3):

\[monthly payment=\frac{ P(\frac{ r }{ M } )}{ 1-(1+\frac{ r }{ M })^{-Mt} }\]

OpenStudy (ac3):

wait how'd you get 349?